# Three-dimensional quantum criticality

### v2

@@ Acknowledgements.
We are grateful to A. Levchenko, N.R. Cooper, T. Helm, T. Hazra, P. Coleman and J. Schmalian for stimulating discussions. We thank S.T. Hannahs for technical advice and assistance. This project was supported by the EPSRC of the UK (grant no. EP/X011992/1). A portion of this work was performed at the National High Magnetic Field Laboratory, which is supported by National Science Foundation Cooperative Agreement Nos. DMR-1644779 & DMR-2128556 and the State of Florida. We acknowledge support of the HLD at HZDR, a member of the European Magnetic Field Laboratory (EMFL). The EMFL also supported dual-access to facilities at MGML, Charles University, Prague, under the European Union’s Horizon 2020 research and innovation programme through the ISABEL project (No. 871106). Crystal growth and characterization were performed in MGML (mgml.eu), which is supported within the program of Czech Research Infrastructures (project no. LM2023065). We acknowledge financial support by the Czech Science Foundation (GACR), project No. 22-22322S. T.I.W. and A.J.H. acknowledge support from EPSRC studentships EP/R513180/1 & EP/M506485/1. T.I.W. and A.G.E. acknowledge support from ICAM through US National Science Foundation (NSF) Grant Number 2201516 under the Accelnet program of the Office of International Science and Engineering and from QuantEmX grants from ICAM and the Gordon and Betty Moore Foundation through Grants GBMF5305 & GBMF9616. D.V.C. acknowledges financial support from the National High Magnetic Field Laboratory through a Dirac Fellowship, which is funded by the National Science Foundation (Grant No. DMR-1644779) and the State of Florida. The work at UW-Madison (D.S.) was financially supported by the National Science Foundation, Quantum Leap Challenge Institute for Hybrid Quantum Architectures and Networks Grant No. OMA-2016136. A.G.E. acknowledges support from the Henry Royce Institute for Advanced Materials through the Equipment Access Scheme enabling access to the Advanced Materials Characterisation Suite at Cambridge, grant numbers EP/P024947/1, EP/M000524/1 & EP/R00661X/1; and from Sidney Sussex College (University of Cambridge).
@@ Appendix A EXPERIMENTAL TECHNIQUES
Contactless electrical conductivity measurements were made by the proximity detector oscillator (PDO) technique [47]. Experiments were performed at the Wuhan National High Magnetic Field Center in pulsed magnetic fields up to 56.7 T, and at the Hochfeld-Magnetlabor Dresden (HLD-EMFL) in a dual-coil magnet to a maximum applied field of 79.7 T. PDO measurements were performed at frequencies of tens of MHz, giving good resolution despite the short duration of a magnetic field pulse, of order 0.1 s (see Figure 15). By tracking the change in frequency $f$ of the PDO circuit [47, 48] as a function of $H$, the change in a sample’s electrical resistivity $\rho$ is captured as
@@ Appendix C SUPPORTING FIGURES
Figure 10: Identification of the spillover of SC3 outside the PPM. (a) DC magnetic field sweeps at $T$ = 0.4 K in the $b-ac$ rotation plane. A subset of these data are presented in Fig. 4a. These data are used to construct the phase diagram in Fig. 4b. (b) DC field sweeps at the elevated temperature of $T$= 6 K – above all critical superconducting temperatures – in order to map the extent of the PPM phase in this rotation plane. The MM transition is observed as a peak in $\rho(H)$ at incremental rotation angles up to $\theta_{b-ac}$ = 27.5°$\degree$; however, at the next measured angle of $\theta_{b-ac}$ = 31.0°$\degree$ no peak is observed up to the highest applied field of 41.5 T. In panel (a) the 0.4 K data at $\theta_{b-ac}$ = 31.0°$\degree$ is plotted in the same blue as panel (b), clearly showing that the zero resistance state of SC3 is accessed even though the MM transition for this field orientation lies at higher field strengths.
Figure 11: Non-Ohmic current–voltage behavior near the QCL. Effective resistivity $\rho$ for incremental currents at isothermal temperature points as indicated. We note that some Joule heating likely occurs for the highest current measurements at 0.5 K (and to a lesser extent for 0.7 K) as seen by the change in slope of $\rho(H)$ in the normal state over the interval 25 T $\lessapprox\mu_{0}H\lessapprox$ 30 T for e.g. the 20 mA curve compared to that at 150 $\upmu$A. However, there is clearly an intrinsic dependence on current for the onset of SC3 between the 150 $\upmu$A and 400 $\upmu$A curves, which exhibit no indication of extrinsic heating effects. Furthermore, at higher temperatures the linear relationship between $V$ and $I$ in the normal state (i.e. at $\mu_{0}H=$ 25 T) holds extremely well for all currents, while even up to 1.8 K at $\mu_{0}H\approx$ 40 T the SC3 onset regime exhibits clearly non-linear $I-V$ behavior. The 1.1 K data are those displayed in Fig. 4e.
@@
Figure 13: Temperature evolution of the metamagnetic transition. PDO data at incremental temperatures for six different magnetic field orientations, as labelled. All data were collected on the same sample, measured in the Wuhan National High Magnetic Field Center.
@@
Figure 16: High field UTe2 phase diagram for magnetic field orientations in the $bc$ and $ba$ rotational planes. The phase boundaries of the SC1, SC2, SC3 and PPM phases as determined by our prior measurements reported in ref. [36] and by those from this study.
Figure 17: Comparison of the angular extent of SC3 in the $b-c$ rotation plane from measurements of electrical conductivity presented in this study along with those reported in refs. [31, 40]. All data are at $T\approx$ 0.5-0.7 K. Orange points represent the transition into the PPM state, while purple and blue symbols indicate the start and end of the SC3 state. Black dashed lines mark the highest angle at which SC3 is still observed; orange dashed lines mark the next subsequent measured angle from each study, where signatures of superconductivity are no longer resolved (only the transition to the PPM phase). Note that two samples are plotted from ref. [40]. The SC3 state seems to terminate in the same range of inclination of 41°$\degree\lessapprox\theta_{b-c}\lessapprox$ 45°$\degree$ in each of these studies.
@@
Figure 19: Comparison of high field PPM and SC3 data measured in this study with those reported in Lewin et al. [32]. Our data are plotted with solid symbols, while those of ref. [32] are given as open symbols. Solid green stars are QCEPs from our measurements, while open green hexagrams are interpreted from the data in ref. [32]. We plot an indicative error bar for the QCEPs interpreted from the data in Lewin et al. [32], as those data were measured below $T_{c}^{SC3}$ therefore it is difficult to definitively determine the location of the QCEP as the domain of SC3 can stretch on either side of the QCL, as we show in Fig. 4. We therefore include these indicative error bars, constructed by spanning the angular region halfway between the last measurement point clearly showing first order–like metamagnetism and the suppression towards 0 K. We find that the dataset reported in Lewin et al. [32] agrees very well with our construction of the UTe2 high magnetic field QCL.

### v3

@@ Acknowledgements.
We are grateful to A. Levchenko, N.R. Cooper, T. Helm, M. Kwasigroch, A. Huxley, T. Hazra, P. Coleman and J. Schmalian for stimulating discussions. We thank S.T. Hannahs for technical advice and assistance. This project was supported by the EPSRC of the UK (grant no. EP/X011992/1). A portion of this work was performed at the National High Magnetic Field Laboratory, which is supported by National Science Foundation Cooperative Agreement Nos. DMR-1644779 & DMR-2128556 and the State of Florida. We acknowledge support of the HLD at HZDR, a member of the European Magnetic Field Laboratory (EMFL). The EMFL also supported dual-access to facilities at MGML, Charles University, Prague, under the European Union’s Horizon 2020 research and innovation programme through the ISABEL project (No. 871106). Crystal growth and characterization were performed in MGML (mgml.eu), which is supported within the program of Czech Research Infrastructures (project no. LM2023065). We acknowledge financial support by the Czech Science Foundation (GACR), project No. 22-22322S. T.I.W. and A.J.H. acknowledge support from EPSRC studentships EP/R513180/1 & EP/M506485/1. T.I.W. and A.G.E. acknowledge support from ICAM through US National Science Foundation (NSF) Grant Number 2201516 under the Accelnet program of the Office of International Science and Engineering and from QuantEmX grants from ICAM and the Gordon and Betty Moore Foundation through Grants GBMF5305 & GBMF9616. D.V.C. acknowledges financial support from the National High Magnetic Field Laboratory through a Dirac Fellowship, which is funded by the National Science Foundation (Grant No. DMR-1644779) and the State of Florida. The work at UW-Madison (D.S.) was financially supported by the National Science Foundation, Quantum Leap Challenge Institute for Hybrid Quantum Architectures and Networks Grant No. OMA-2016136. A.G.E. acknowledges support from the Henry Royce Institute for Advanced Materials through the Equipment Access Scheme enabling access to the Advanced Materials Characterisation Suite at Cambridge, grant numbers EP/P024947/1, EP/M000524/1 & EP/R00661X/1; and from Sidney Sussex College (University of Cambridge).
@@ Appendix A EXPERIMENTAL TECHNIQUES
Contactless electrical conductivity measurements were made by the proximity detector oscillator (PDO) technique [73], following the same methodology as our prior UTe2 PDO study [74]. Experiments were performed at the Wuhan National High Magnetic Field Center in pulsed magnetic fields up to 56.7 T, and at the Hochfeld-Magnetlabor Dresden (HLD-EMFL) in a dual-coil magnet to a maximum applied field of 79.7 T. PDO measurements were performed at frequencies of tens of MHz, giving good resolution despite the short duration of a magnetic field pulse, of order 0.1 s (see Figure 17). By tracking the change in frequency $f$ of the PDO circuit [73, 75] as a function of $H$, the change in a sample’s electrical resistivity $\rho$ is captured as
@@
Figure 11: Schematic representation of the experimental setup used for simultaneously applying $H_{a}$, $H_{b}$ and $H_{c}$ magnetic field components to UTe2. Samples were mounted on G10 wedges, giving a fixed inclination between the $c$-axis and either $a$ or $b$. By rotating the orientation of the magnetic field about an axis orthogonal to the axial inclination of the wedge, the applied components of $H_{a}$, $H_{b}$ and $H_{c}$ were modulated.
@@ Appendix C SUPPORTING FIGURES
Figure 12: Identification of the spillover of SC3 outside the PPM. (a) DC magnetic field sweeps at $T$ = 0.4 K in the $b-ac$ rotation plane with $J=$ 150 mA/cm2. A subset of these data are presented in Fig. 5a. These data are used to construct the phase diagram in Fig. 5b. (b) DC field sweeps at the elevated temperature of $T$= 6 K – above all critical superconducting temperatures – in order to map the extent of the PPM phase in this rotation plane. The MM transition is observed as a peak in $\rho(H)$ at incremental rotation angles up to $\theta_{b-ac}$ = 27.5°$\degree$; however, at the next measured angle of $\theta_{b-ac}$ = 31.0°$\degree$ no peak is observed up to the highest applied field of 41.5 T. In panel (a) the 0.4 K data at $\theta_{b-ac}$ = 31.0°$\degree$ is plotted in the same blue as panel (b), clearly showing that the zero resistance state of SC3 is accessed even though the MM transition for this field orientation lies at higher field strengths.
Figure 13: Non-Ohmic current–voltage behavior near the QCL. Effective resistivity $\rho$ for incremental currents at isothermal temperature points as indicated. We note that some Joule heating likely occurs for the highest current measurements at 0.5 K (and to a lesser extent for 0.7 K) as seen by the change in slope of $\rho(H)$ in the normal state over the interval 25 T $\lessapprox\mu_{0}H\lessapprox$ 30 T for e.g. the 2.5 A/cm2 curve compared to that at 50 mA/cm2. However, there is clearly an intrinsic dependence on current for the onset of SC3 between the 150 mA/cm2 and 500 mA/cm2 curves, which exhibit no indication of extrinsic heating effects. Furthermore, at higher temperatures the linear relationship between $V$ and $J$ in the normal state (i.e. at $\mu_{0}H=$ 25 T) holds extremely well for all currents, while even up to 1.8 K at $\mu_{0}H\approx$ 40 T the SC3 onset regime exhibits clearly non-linear $J-V$ behavior. The 1.1 K data are those displayed in Fig. 5e.
@@
Figure 15: Temperature evolution of the metamagnetic transition. PDO data at incremental temperatures for six different magnetic field orientations, as labeled. All data were collected on the same sample, measured in the Wuhan National High Magnetic Field Center.
@@
Figure 18: High field UTe2 phase diagram for magnetic field orientations in the $bc$ and $ba$ rotational planes. The phase boundaries of the SC1, SC2, SC3 and PPM phases are determined by our prior measurements reported in refs. [25, 50] and by those from this study.
Figure 19: Comparison of the angular extent of SC3 in the $b-c$ rotation plane from measurements of electrical conductivity presented in this study along with those reported in refs. [33, 49]. All data are at $T\approx$ 0.5-0.7 K. Orange points represent the transition into the PPM state, while purple and blue symbols indicate the start and end of the SC3 state. Black dashed lines mark the highest angle at which SC3 is still observed; orange dashed lines mark the next subsequent measured angle from each study, where signatures of superconductivity are no longer resolved (only the transition to the PPM phase). Note that two samples are plotted from ref. [49]. The SC3 state seems to terminate over a similar range of inclination of 41°$\degree\lessapprox\theta_{b-c}\lessapprox$ 45°$\degree$ in each of these studies.
@@
Figure 21: Comparison of high field PPM and SC3 data measured in this study with those reported in Lewin et al. [36]. Our data are plotted with solid symbols, while those of ref. [36] are given as open symbols. Solid green stars are QCEPs from our measurements, while open green hexagrams are interpreted from the data in ref. [36]. We plot an indicative error bar for the QCEPs interpreted from the data in Lewin et al. [36], as those data were measured below $T_{c}^{SC3}$ therefore it is difficult to definitively determine the location of the QCEP as the domain of SC3 can stretch on either side of the QCL, as we show in Fig. 5. We therefore include these indicative error bars, constructed by spanning the angular region halfway between the last measurement point clearly showing first-order-like metamagnetism and the suppression towards 0 K. We find that the dataset reported in Lewin et al. [36] agrees very well with our construction of the UTe2 high magnetic field QCL.